# Environment-Driven Online LiDAR-Camera Extrinsic Calibration

### v2

@@ Abstract
LiDAR-camera extrinsic calibration (LCEC) is crucial for multi-modal data fusion in mechatronics. Existing methods, whether target-based or target-free, typically rely on customized calibration targets or fixed scene types, limiting their practicality in real-world applications. To address these challenges, we introduce EdO-LCEC, the first environment-driven online calibration approach. Unlike traditional target-free methods, EdO-LCEC observes the feature density of the application environment through a generalizable scene discriminator. Based on this feature density, EdO-LCEC extracts LiDAR intensity and depth features from varying perspectives to achieve higher calibration accuracy. To overcome the challenges of cross-modal feature matching between LiDAR and camera, we propose dual-path correspondence matching (DPCM), which leverages both structural and textural consistency for reliable 3D-2D correspondences. Additionally, our approach models the calibration process as a joint optimization problem utilizing global constraints from multiple views and scenes to enhance accuracy. Extensive experiments on real-world datasets demonstrate that EdO-LCEC outperforms state-of-the-art methods, particularly in sparse or partially overlapping sensor views.
Index Terms:
Environment-driven, LiDAR-camera extrinsic calibration, multi-modal data fusion.
[…]
figure Visualization of calibration results through LiDAR and camera data fusion in KITTI odometry 00 sequence: (a)-(i) zoomed-in regions that illustrate the alignment between the camera images and the LiDAR point clouds.
@@ I Introduction
Perception is a critical challenge for autonomous mobile robots [1]. As the front-end module of mechatronic systems, it provides essential environmental information for high-level navigation. Modern LiDAR-camera fusion systems significantly enhance robotic perception capabilities [2, 3, 4, 5]. While LiDARs provide accurate spatial information, cameras capture rich textural details [6]. Their complementary fusion enables robust performance in key robotic tasks including SLAM [7, 8] and object recognition [9]. LiDAR-camera extrinsic calibration (LCEC), which estimates the relative pose between the two sensors, becomes a core and foundational process for effective data fusion in mechatronics.
@@
Current LiDAR-camera extrinsic calibration methods are primarily categorized as either target-based or target-free. Target-based approaches [10, 11, 12, 13, 14, 15] have long been the preferred choice in this field. They are typically offline, relying on customized calibration targets (typically checkerboards). However, they often demonstrate poor adaptability to real-world environments. This is largely because extrinsic parameters may change significantly due to moderate shocks or during extended operations in environments with vibrations. Online, target-free approaches aim at overcoming this problem by extracting informative visual features directly from the environment. Previous works [16, 17] estimate the extrinsic parameters by matching the cross-modal edges between LiDAR projections and camera images. While effective in specific scenarios with abundant features, these traditional methods heavily rely on well-distributed edge features. Recent advances in deep learning techniques have spurred extensive explorations, such as [18, 19, 20], to leverage semantic information to aid cross-modal feature matching. Although these approaches have shown compelling performance in specific scenarios, such as urban freeway, they predominantly rely on curated, pre-defined objects, e.g., vehicles, lanes, and traffic poles. On the other hand, several end-to-end deep learning networks [21, 22, 23, 24, 25, 26] have been developed to find a more direct solution for LCEC. While these methods have demonstrated effectiveness on their training datasets like KITTI [27], they highly rely on the training setup and are thus less generalizable. In our previous work, MIAS-LCEC [6], we developed a two-stage coarse-to-fine matching approach using the large vision model MobileSAM [28] to improve cross-modal feature matching. While this method achieves high accuracy with solid-state LiDARs (producing dense point clouds) when the sensors have overlapping fields of view, its performance significantly degrades in challenging scenarios with sparse or incomplete point clouds.
@@
Although target-free LCEC methods eliminate the reliance on calibration targets, they struggle with complex, unstructured scenes due to their inability to adapt to environmental changes. To address this, we introduce environmental observation into target-free calibration, proposing the first environment-driven framework, EdO-LCEC. EdO-LCEC adapts to external conditions to maintain optimal performance by balancing the feature density between LiDAR projection and camera image. In this way, the system is no longer constrained by fixed strategies but can intelligently respond to environmental changes. Specifically, as illustrated in Fig. Environment-Driven Online LiDAR-Camera Extrinsic Calibration, we consider the working environment of the sensors as a sequence composed of multiple scenes. By selectively merging multiple scenes across different times, this approach could achieve high-precision calibration dynamically. A prerequisite for the success of EdO-LCEC is the design of a generalizable scene discriminator. The scene discriminator employs large vision models to conduct depth estimation and image segmentation. In detail, it calculates the feature density of the calibration scene and uses it to guide the generation of multiple virtual cameras for projecting LiDAR intensities and depth. This improved LiDAR point cloud projecting strategy increases the available environmental features, and thus overcomes the previous reliance of algorithms [16, 29] on uniformly distributed geometric or textural features. At each scene, we perform dual-path correspondence matching (DPCM). Different from the C3M in MIAS-LCEC [6], DPCM divides correspondence matching into spatial and textural pathways to fully leverage both geometric and semantic information. Each pathway constructs a cost matrix based on structural and textural consistency, guided by accurate semantic priors, to yield reliable 3D-2D correspondences. Finally, the correspondences obtained from multiple views and scenes are used as inputs for our proposed multi-view and multi-scene joint optimization, which derives and refines the extrinsic matrix between LiDAR and camera. Through extensive experiments conducted on three real-world datasets, EdO-LCEC demonstrates superior robustness and accuracy compared to other SoTA approaches.
@@
Generalizable scene discriminator, which can automatically observe the calibration scene by evaluating the feature density through potential spatial, textural and semantic features extracted by SoTA LVMs.
@@ II Related Work
Advances in deep learning techniques have driven significant exploration into enhancing traditional target-free algorithms. Some studies [37, 38, 18, 19, 20, 39, 29, 6] explore attaching deep learning modules to their calibration framework as useful tools to enhance calibration efficiency. For instance, [38] accomplishes LiDAR and camera registration by aligning road lanes and poles detected by semantic segmentation. Similarly, [19] employs stop signs as calibration primitives and refines results over time using a Kalman filter. A recent study [29] introduced Direct Visual LiDAR Calibration (DVL), a novel point-based method that utilizes SuperGlue [40] to establish direct 3D-2D correspondences between LiDAR and camera data. On the other hand, several learning-based algorithms [21, 22, 23, 24, 25, 26] try to convert the calibration process into more direct approaches by exploiting end-to-end deep learning networks. Although these methods have demonstrated effectiveness on public datasets like KITTI [27], which primarily focuses on urban driving scenarios, their performance has not been extensively validated on other types of real-world datasets that contain more challenging scenes. Furthermore, their dependence on pre-defined sensor configurations (both LiDAR and camera) poses implementation challenges. In comparison, our approach introduces feature density evaluation to achieve environmental observation. This environment-driven strategy enables robots to maintain high-precision calibration anytime and anywhere.
@@ III Methodology
where represents the rotation matrix, denotes the translation vector, and represents a column vector of zeros. We first give an overview of the proposed method, as shown in Fig. 1. It mainly contains three stages:
@@ III-A Generalizable Scene Discriminator
where denotes the textural density and represents the structural density. The occlusion challenges caused by the different perspectives of LiDAR and the camera [16, 6], combined with limited feature availability in sparse point clouds, mean that a single pair of virtual cameras is insufficient for a comprehensive view of the calibration scene. Let represent the event of capturing enough effective features, with probability . If we have intensity virtual cameras and depth virtual cameras, the probability of capturing enough effective features is . In theory, as , the probability , leading to . Increasing the number of virtual cameras raises the likelihood of capturing more potential correspondences, thus enhancing calibration accuracy. However, applying an infinite number of virtual cameras during calibration is impractical. Considering the trade-off between calibration accuracy and computational resources, we set the number of multiple virtual cameras to satisfy the balance of feature density:
As depicted in Fig. 2, in practical applications, we set multiple virtual cameras inside a sphere originating from the LiDAR perspective center. Since the feature density is similar if the perspective viewpoints of virtual cameras are close to the initial position, we can assume that and . So and can be obtained as follows:
@@ III-B Dual-Path Correspondence Matching
Given the segmented masks with detected corner points, dual-path correspondence matching leverages them to achieve dense and reliable 3D-2D correspondences. DPCM consists of two pathways, one for correspondence matching of LIP and RGB images, and the other for LDP and depth images. For each pathway, DPCM adopted the approach outlined in [6] to obtain mask matching result from a cost matrix . Each matched mask pair can estimate an affine transformation to guide the correspondence matching. Specifically, we update the corner points in the virtual image to a location that is close to its true projection coordinate using this affine transformation, as follows:
@@
denotes the matching cost between the -th corner point of a mask in the LiDAR virtual image and the -th corner point of a mask in the camera image. (6) consists of structural and textural consistency. As illustrated in Fig. 3, the structural consistency measures the structural difference of corner points in the virtual and real image, where denotes the average perimeter of the matched masks and represents the similarity of the neighboring vertices between current and target corner point. The textural consistency derives from the relative textural similarity of the neighboring zone. After establishing the cost matrix, a strict criterion is applied to achieve reliable matching. Matches with the lowest costs in both horizontal and vertical directions of are determined as the optimum corner point matches. Since every can trace back to a LiDAR 3D point , and every is related to a pixel (represented in homogeneous coordinates as ) in the camera image, the final correspondence matching result of DPCM is .
@@ III-C Multi-View and Multi-Scene Joint Optimization
where is the correspondence subset in the -th scene with the -th smallest mean reprojection error under its PnP solution. denotes the number of correspondences in the -th scene. and represent the maximum number of subsets across all scenarios and the maximum number of reliable subsets in a single scene, respectively. The multi-scene optimization process then solves the final extrinsic matrix by minimizing the joint reprojection error:
across multiple spaces in the environment at different times. This process combines optimal correspondences from both spatial and textural pathways in each scenario, enabling robust environment-driven optimization through multi-view and multi-scene fusion. As a result, our method achieves human-like adaptability to dynamic environments. This environment perception make EdO-LCEC behave much better than our previous method MIAS-LCEC, especially in conditions when point clouds are sparse or incomplete.
@@ IV-A Experimental Setup and Evaluation Metrics
| CRLF [38] | - | 0.629 | 4.118 | 0.033 | 0.464 | 0.416 | 3.648 | 1.483 | 0.550 | 0.633 | 4.606 | 0.039 | 0.458 | 0.424 | 4.055 | 1.636 | 0.644 |
| UMich [17] | - | 4.161 | 0.321 | 0.113 | 3.111 | 2.138 | 0.286 | 0.077 | 0.086 | 4.285 | 0.329 | 0.108 | 3.277 | 2.088 | 0.290 | 0.085 | 0.090 |
| HKU-Mars [16] | - | 33.84 | 6.355 | 19.89 | 18.71 | 19.32 | 3.353 | 3.232 | 2.419 | 32.89 | 4.913 | 18.99 | 15.77 | 17.00 | 2.917 | 2.564 | 1.646 |
@@
| MIAS-LCEC [6] | - | 5.385 | 1.014 | 1.574 | 4.029 | 4.338 | 0.724 | 0.383 | 0.343 | 7.655 | 1.342 | 1.910 | 5.666 | 6.154 | 0.843 | 0.730 | 0.358 |
| EdO-LCEC (Ours) | - | 0.295 | 0.082 | 0.117 | 0.176 | 0.150 | 0.051 | 0.044 | 0.032 | 0.336 | 0.118 | 0.216 | 0.168 | 0.121 | 0.083 | 0.067 | 0.032 |
@@
| CRLF [38] | 0.623 | 7.363 | 0.632 | 3.642 | 0.845 | 6.007 | 0.601 | 0.372 | 0.616 | 5.961 | 0.615 | 25.762 | 0.606 | 1.807 | 0.625 | 5.376 | 0.626 | 5.133 |
| UMich [17] | 2.196 | 0.305 | 3.733 | 0.331 | 3.201 | 0.316 | 2.086 | 0.348 | 3.526 | 0.356 | 2.914 | 0.353 | 3.928 | 0.368 | 3.722 | 0.367 | 3.117 | 0.363 |
| HKU-Mars [16] | 20.73 | 3.770 | 32.95 | 12.70 | 21.99 | 3.493 | 4.943 | 0.965 | 34.42 | 6.505 | 25.20 | 7.437 | 33.10 | 7.339 | 26.62 | 8.767 | 20.38 | 3.459 |
@@
| MIAS-LCEC [6] | 0.621 | 0.300 | 0.801 | 0.327 | 1.140 | 0.324 | 0.816 | 0.369 | 4.768 | 0.775 | 2.685 | 0.534 | 11.80 | 1.344 | 5.220 | 0.806 | 0.998 | 0.432 |
| EdO-LCEC (Ours) | 2.269 | 0.459 | 0.561 | 0.142 | 0.737 | 0.137 | 1.104 | 0.339 | 0.280 | 0.093 | 0.485 | 0.124 | 0.188 | 0.076 | 0.352 | 0.115 | 0.386 | 0.120 |
| Approach | Residential Community | Urban Freeway | Building | Challenging Weather | Indoor | Challenging Illumination | All | |||||||
| (∘) | (m) | (∘) | (m) | (∘) | (m) | (∘) | (m) | (∘) | (m) | (∘) | (m) | (∘) | (m) | |
| CRLF [38] | 1.594 | 0.464 | 1.582 | 0.140 | 1.499 | 20.17 | 1.646 | 2.055 | 1.886 | 30.05 | 1.876 | 19.05 | 1.683 | 11.13 |
| UMich [17] | 4.829 | 0.387 | 2.267 | 0.166 | 11.914 | 0.781 | 1.851 | 0.310 | 2.029 | 0.109 | 5.012 | 0.330 | 4.265 | 0.333 |
| HKU-Mars [16] | 2.695 | 1.208 | 2.399 | 1.956 | 1.814 | 0.706 | 2.578 | 1.086 | 2.527 | 0.246 | 14.996 | 3.386 | 3.941 | 1.261 |
| DVL [29] | 0.193 | 0.063 | 0.298 | 0.124 | 0.200 | 0.087 | 0.181 | 0.052 | 0.391 | 0.030 | 1.747 | 0.377 | 0.423 | 0.100 |
| MIAS-LCEC [6] | 0.190 | 0.050 | 0.291 | 0.111 | 0.198 | 0.072 | 0.177 | 0.046 | 0.363 | 0.024 | 0.749 | 0.118 | 0.298 | 0.061 |
| EdO-LCEC (Ours) | 0.168 | 0.044 | 0.293 | 0.105 | 0.184 | 0.057 | 0.183 | 0.044 | 0.338 | 0.027 | 0.474 | 0.104 | 0.255 | 0.055 |
In our experiments, we compare our proposed EdO-LCEC with SoTA online, target-free LCEC approaches on the public dataset KITTI odometry [27] (including 00-09 sequences) and MIAS-LCEC [6] (including target-free datasets MIAS-LCEC-TF70 and MIAS-LCEC-TF360). To comprehensively evaluate the performance of LCEC approaches, we use the magnitude of Euler angle error and the magnitude 222The translation from LiDAR pose to camera pose is when (1) is used to depict the point translation. of the translation error, with the following expression:
to quantify the calibration errors. In (11), and represent the estimated and ground-truth Euler angle vectors, computed from the rotation matrices and , respectively. Similarly, and denote the estimated and ground-truth translation vectors from LiDAR to camera, respectively.
Notably, sequences in KITTI odometry, aside from 00, were included in the training datasets for the learning-based methods [22, 26, 24]. To ensure a fair comparison, we reproduced calibration results for both the left and right cameras on sequence 00 when the authors provided their code; otherwise, we used the reported results for the left camera from their papers. Since most learning-based methods lack APIs for custom data, our comparison with these methods is limited to the KITTI odometry 00 sequence. For experiments on the MIAS-LCEC dataset, as the results for the compared methods [38, 16, 29, 6, 17] are reported in [6], we directly use the values presented in that paper.
@@
IV-B Comparison with State-of-the-Art Method
In this section, quantitative comparisons with SoTA approaches on three datasets are presented in Fig. 7, Tables I333The results generated by LCCNet using the authors’ code yield higher errors compared to those reported in their paper., II, III, and IV. Additionally, qualitative results are illustrated in Figs. 4 and 5.
IV-B1 Evaluation on KITTI Odometry
The results shown in Table I and II suggest that, with the exception of sequences 01 and 04, our method achieves SoTA performance across the ten sequences (00-09) in KITTI odometry. Specifically, in the 00 sequence, EdO-LCEC reduces the by around 35.2-99.8% and the by 11.8-98.7% for the left camera, and reduces the by around 46.9-99.7% and the by 13.9-97.6% for the right camera. Additionally, according to Fig. 5, it can be observed that the point cloud of a single frame in KITTI is so sparse that the other approaches behave poorly. In contrast, our proposed method overcomes this difficulty and achieves high-quality data fusion through the calibration result. We attribute these performance improvements to our multi-view and multi-scene joint optimization. Merging the optimal matching results from multiple views and scenes maximizes the number of reliable correspondences and ultimately improves overall calibration accuracy.
IV-B2 Evaluation on MIAS-LCEC Dataset

### v3

@@ Abstract
LiDAR-camera extrinsic calibration (LCEC) is crucial for multi-modal data fusion in autonomous robotic systems. Existing methods, whether target-based or target-free, typically rely on customized calibration targets or fixed scene types, which limit their applicability in real-world scenarios. To address these challenges, we present EdO-LCEC, the first environment-driven online calibration approach. Unlike traditional target-free methods, EdO-LCEC employs a generalizable scene discriminator to estimate the feature density of the application environment. Guided by this feature density, EdO-LCEC extracts LiDAR intensity and depth features from varying perspectives to achieve higher calibration accuracy. To overcome the challenges of cross-modal feature matching between LiDAR and camera, we introduce dual-path correspondence matching (DPCM), which leverages both structural and textural consistency for reliable 3D-2D correspondences. Furthermore, we formulate the calibration process as a joint optimization problem that integrates global constraints across multiple views and scenes, thereby enhancing overall accuracy. Extensive experiments on real-world datasets demonstrate that EdO-LCEC outperforms state-of-the-art methods, particularly in scenarios involving sparse point clouds or partially overlapping sensor views.
Note to Practitioners
This article presents an environment-driven approach for LiDAR-camera extrinsic calibration. Unlike conventional target-free methods, the proposed EdO-LCEC not only extracts matchable features from real-world scenes but also adapts its calibration strategy based on the environmental feature density. This environmental awareness significantly enhances calibration robustness. By focusing on cross-modal feature matching and extrinsic optimization, our method performs reliably across various sensor configurations, including solid-state and mechanical LiDARs with differing fields of view and point densities. The proposed approach offers a practical and generalizable solution that improves upon existing target-free methods, facilitating deployment in sensor fusion and mechatronic systems. Our calibration software developed on EdO-LCEC will be publicly available at https://mias.group/EdO-LCEC.
@@ I-A Background
Perception is a fundamental capability in autonomous mobile robotics, but achieving robust and generalizable perception remains a significant challenge [1]. By allowing robots to acquire and interpret information about their surroundings, it provides the critical foundation for dependable navigation, planning, and high-level decision-making [2]. Modern LiDAR-camera fusion systems significantly enhance robotic perception capabilities [3]. While LiDARs provide accurate spatial information, cameras capture rich textural details [4, 5]. Their complementary fusion enables robust performance in key robotic tasks, including SLAM [6, 7, 8, 9, 10], object recognition [11, 12], and localization [13]. LiDAR-camera extrinsic calibration (LCEC), which estimates the relative pose between the two sensors, becomes a core and foundational process for effective data fusion in automation science and engineering. Extensive research on offline, target-based LCEC has yielded numerous effective and robust algorithms over the decades. However, online target-free methods still struggle in complex, unstructured scenes due to limited adaptability to the working environment, and much work remains to fully realize their potential.
I-B Existing Challenges and Motivation
Current LCEC methods are primarily categorized as either target-based or target-free. Target-based approaches [14, 15, 16, 17, 18] have long been the preferred choice in this field. They are typically offline, relying on customized calibration targets (typically checkerboards). However, they often demonstrate poor adaptability to real-world environments. This is largely because extrinsic parameters may change significantly due to moderate shocks or during extended operations in environments with vibrations.
@@ I-B Existing Challenges and Motivation
Online, target-free approaches aim at overcoming this problem by extracting informative visual features directly from the environment. Previous works [19, 20] estimate the extrinsic parameters by matching the cross-modal edges between LiDAR projections and camera images. While effective in specific scenarios with abundant features, these traditional methods heavily rely on well-distributed edge features. Recent advances in deep learning techniques have spurred extensive explorations, such as [21, 22, 23], to leverage semantic information to aid cross-modal feature matching. Although these approaches have shown compelling performance in specific scenarios, such as urban freeway, they predominantly rely on curated, pre-defined objects, e.g., vehicles, lanes, and traffic poles. On the other hand, several end-to-end deep learning networks [24, 25, 26, 27, 28, 29] have been developed to find a more direct solution for LCEC. While these methods have demonstrated effectiveness on their training datasets, like KITTI [30], they highly rely on the training setup and are thus less generalizable. The recent work MIAS-LCEC [4] exploits the large vision model MobileSAM [31] to improve cross-modal feature matching. While this method achieves high accuracy with solid-state LiDARs (producing dense point clouds) when the sensors have overlapping fields of view, its performance significantly degrades in challenging scenarios with sparse or incomplete point clouds.
Although current target-free LCEC methods eliminate reliance on calibration targets, they are still restricted to specific sensor types and pre-defined environmental settings. As a result, they are hard to balance calibration accuracy and generalizability in complex, unstructured scenes due to limited adaptability. This challenge becomes even more pronounced in dynamic scenarios where sensor poses are uncertain and frequently changing. Therefore, this study aims to create a more flexible online calibration framework that adapts its strategy based on the environmental cues, thereby improving the robustness of online, target-free LCEC.
I-C Novel Contributions
@@
In this article, we move one step forward to introduce environmental observation into target-free calibration, proposing the first environment-driven framework, EdO-LCEC. EdO-LCEC adapts to external conditions to maintain optimal performance by balancing the feature density between LiDAR projection and camera image. Unlike the conventional target-free approaches, EdO-LCEC is no longer constrained by fixed strategies but can intelligently respond to environmental changes. Specifically, as illustrated in Fig. 1, we consider the working environment of the sensors as a sequence composed of multiple scenes. By actively perceiving the feature density of the environment and merging multiple scenes across different times, this approach could achieve high-precision calibration dynamically. A prerequisite for the success of EdO-LCEC is the design of a generalizable scene discriminator. The scene discriminator employs large vision models to conduct depth estimation and image segmentation. In detail, it calculates the feature density of the calibration scene and uses it to guide the generation of multiple virtual cameras for projecting LiDAR intensities and depth. This improved LiDAR point cloud projecting strategy increases the available environmental features, and thus overcomes the previous reliance of algorithms [19, 32] on uniformly distributed geometric or textural features. At each scene, we perform dual-path correspondence matching (DPCM). Different from the C3M in MIAS-LCEC [4], DPCM divides correspondence matching into spatial and textural pathways to fully leverage both geometric and semantic information. Each pathway constructs a cost matrix based on structural and textural consistency, guided by accurate semantic priors, to yield reliable 3D-2D correspondences. Finally, the correspondences obtained from multiple views and scenes are used as inputs for our proposed multi-view and multi-scene joint optimization, which derives and refines the extrinsic matrix between LiDAR and camera. Through extensive experiments conducted on three real-world datasets, EdO-LCEC demonstrates superior robustness and accuracy compared to other SoTA approaches.
@@
Generalizable scene discriminator, which can automatically observe the calibration scene by evaluating the feature density through potential spatial, textural, and semantic features extracted by SoTA LVMs.
@@ II Related Work
Advances in deep learning techniques have driven significant exploration into enhancing traditional target-free algorithms. Some studies [40, 21, 22, 32, 4] explore attaching deep learning modules to their calibration framework as useful tools to enhance calibration efficiency. For instance, [40] accomplishes LiDAR and camera registration by aligning road lanes and poles detected by semantic segmentation. Similarly, [22] employs stop signs as calibration primitives and refines results over time using a Kalman filter. A recent study [32] introduced Direct Visual LiDAR Calibration (DVL), a novel point-based method that utilizes SuperGlue [41] to establish direct 3D-2D correspondences between LiDAR and camera data. On the other hand, several learning-based algorithms [25, 26, 27, 28, 29] attempt to reformulate the calibration process into more direct solutions by leveraging end-to-end deep learning networks. Although these methods have shown promising results on public datasets such as KITTI [30], which primarily focus on urban driving scenarios, their performance has not been extensively validated on other real-world datasets that include more diverse and challenging scenes. Moreover, end-to-end models trained on a single LiDAR-camera pair often overfit to the intrinsic and extrinsic parameters of the training setup and fail to generalize, even when evaluated on a different camera within the same dataset. For instance, a network trained end-to-end on the left camera of KITTI but evaluated on the right camera still predicts the pose of the left camera. This indicates that such networks primarily memorize specific extrinsic parameters rather than learning a generalizable calibration strategy. In contrast, our method incorporates feature density evaluation to achieve environment-aware calibration. This environment-driven strategy enables robots to maintain high-precision calibration across diverse conditions without being constrained by specific sensor configurations.
@@ III Methodology
where represents the rotation matrix, denotes the translation vector, and represents a column vector of zeros. We first give an overview of the proposed method. As shown in Fig. 2, it mainly contains three stages:
@@ III-A Generalizable Scene Discriminator
where denotes the textural density and represents the structural density. The occlusion challenges caused by the different perspectives of LiDAR and the camera [19, 4], combined with limited feature availability in sparse point clouds, mean that a single pair of virtual cameras is insufficient for a comprehensive view of the calibration scene. Let represent the event of capturing enough effective features, with probability . If we have intensity virtual cameras and depth virtual cameras, the probability of capturing enough effective features is . In theory, as , the probability , leading to . Increasing the number of virtual cameras raises the likelihood of capturing more potential correspondences, thus enhancing calibration accuracy.
However, employing an infinite number of virtual cameras during calibration is impractical. In real applications, a straightforward approach is to let users specify the number of virtual cameras based on their experience. If an initial estimate of the calibration parameters is available, the number of virtual cameras can also be inferred from the distance between the LiDAR origin and the initial estimate. Another automated option is to compute the ratio between the LiDAR and camera fields of view (FoVs) to determine the minimum number of virtual cameras required. Each virtual camera replicates the intrinsic FoV of the real camera, and together they fully cover the LiDAR scanning range. While the above strategies enhanced calibration performance in certain specific cases, they lacked generalizability across more complex and diverse scenarios. The required parameters vary with environmental changes, and the absence of environmental perception prevents the algorithm from automatically adapting to unseen and challenging conditions.
In this work, we introduce a fully automated strategy for generating virtual cameras guided by environmental information. Considering the trade-off between calibration accuracy and computational cost, we determine the number of virtual cameras to balance the feature density:
As depicted in Fig. 3, in practical applications, we set multiple virtual cameras inside a sphere originating from the LiDAR perspective center. A smaller or larger sphere might be used to set more cameras if necessary. Since the feature density is similar if the perspective viewpoints of virtual cameras are close to the initial position, we can assume that and . So and can be obtained as follows:
@@ III-B Dual-Path Correspondence Matching
Given the segmented masks with detected corner points, dual-path correspondence matching leverages them to achieve dense and reliable 3D-2D correspondences. DPCM consists of two pathways, one for correspondence matching of LIP and RGB images, and the other for LDP and depth images. For each pathway, DPCM adopted the approach outlined in [4] to obtain mask matching result from a cost matrix . Each matched mask pair can estimate a 4-DoF similarity transform to guide the correspondence matching. Specifically, we update the corner points in the virtual image to a location that is close to its true projection coordinate using this affine transformation, as follows:
@@
denotes the matching cost between the -th corner point of a mask in the LiDAR virtual image and the -th corner point of a mask in the camera image. (6) consists of structural and textural consistency. As illustrated in Fig. 4, the structural consistency measures the structural difference of corner points in the virtual and real image, where serves as a width parameter based on the average perimeter of the matched masks and represents the similarity of the neighboring vertices between current and target corner point. The textural consistency derives from the relative textural similarity of the neighboring zone. After establishing the cost matrix, a strict criterion is applied to achieve reliable matching. Matches with the lowest costs in both horizontal and vertical directions of are determined as the optimum corner point matches. Since every can trace back to a LiDAR 3D point , and every is related to a pixel (represented in homogeneous coordinates as ) in the camera image, the final correspondence matching result of DPCM is .
@@ III-C Multi-View and Multi-Scene Joint Optimization
where is the -th selected correspondences subset in the -th scene. denotes the number of correspondences in the -th scene. represents the maximum number of correspondences hope to get from all scenarios, and denotes the maximum number of reliable correspondences in a single scene. The multi-scene optimization process then solves the final extrinsic matrix by minimizing the joint reprojection error:
across multiple spaces in the environment at different times. This process combines optimal correspondences from both spatial and textural pathways in each scenario, enabling robust environment-driven optimization through multi-view and multi-scene fusion. As a result, our method achieves human-like adaptability to dynamic environments. This environment perception makes EdO-LCEC behave much better than other target-free approaches, especially in conditions when point clouds are sparse or incomplete.
@@ IV-A Experimental Setup and Evaluation Metrics
| CRLF [40] | - | 0.629 | 4.116 | 0.033 | 0.464 | 0.416 | 3.648 | 1.473 | 0.550 | 0.633 | 4.606 | 0.039 | 0.458 | 0.424 | 4.055 | 1.636 | 0.644 |
| UMich [20] | - | 4.161 | 0.319 | 0.113 | 3.111 | 2.138 | 0.286 | 0.068 | 0.086 | 4.285 | 0.329 | 0.108 | 3.277 | 2.088 | 0.290 | 0.085 | 0.090 |
| HKU-Mars [19] | - | 33.84 | 6.354 | 19.89 | 18.71 | 19.32 | 3.353 | 3.228 | 2.419 | 32.89 | 4.913 | 18.99 | 15.77 | 17.00 | 2.917 | 2.564 | 1.646 |
@@
| MIAS-LCEC [4] | - | 5.385 | 1.013 | 1.574 | 4.029 | 4.338 | 0.724 | 0.373 | 0.343 | 7.655 | 1.342 | 1.910 | 5.666 | 6.154 | 0.843 | 0.730 | 0.358 |
| EdO-LCEC (Ours) | - | 0.295 | 0.078 | 0.117 | 0.176 | 0.150 | 0.051 | 0.038 | 0.032 | 0.336 | 0.118 | 0.216 | 0.168 | 0.121 | 0.083 | 0.067 | 0.032 |
@@
| CRLF [40] | 0.623 | 7.363 | 0.632 | 3.640 | 0.845 | 6.007 | 0.601 | 0.372 | 0.616 | 5.961 | 0.615 | 25.762 | 0.606 | 1.807 | 0.625 | 5.376 | 0.626 | 5.133 |
| UMich [20] | 2.196 | 0.303 | 3.733 | 0.329 | 3.201 | 0.316 | 2.086 | 0.348 | 3.526 | 0.356 | 2.914 | 0.353 | 3.928 | 0.368 | 3.722 | 0.367 | 3.117 | 0.363 |
| HKU-Mars [19] | 20.73 | 3.768 | 32.95 | 12.69 | 21.99 | 3.493 | 4.943 | 0.965 | 34.42 | 6.505 | 25.20 | 7.437 | 33.10 | 7.339 | 26.62 | 8.767 | 20.38 | 3.459 |
@@
| MIAS-LCEC [4] | 0.621 | 0.298 | 0.801 | 0.324 | 1.140 | 0.324 | 0.816 | 0.369 | 4.768 | 0.775 | 2.685 | 0.534 | 11.80 | 1.344 | 5.220 | 0.806 | 0.998 | 0.432 |
| EdO-LCEC (Ours) | 2.269 | 0.462 | 0.561 | 0.137 | 0.737 | 0.137 | 1.104 | 0.339 | 0.280 | 0.093 | 0.485 | 0.124 | 0.188 | 0.076 | 0.352 | 0.115 | 0.386 | 0.120 |
| Approach |
Urban Freeway | Building |
Indoor |
All | |||||||||||||
| (∘) | (m) | (∘) | (m) | (∘) | (m) | (∘) | (m) | (∘) | (m) | (∘) | (m) | (∘) | (m) | |||||||
| CRLF [40] | 1.594 | 0.464 | 1.582 | 0.140 | 1.499 | 20.17 | 1.646 | 2.055 | 1.886 | 30.05 | 1.876 | 19.05 | 1.683 | 11.13 | ||||||
| UMich [20] | 4.829 | 0.387 | 2.267 | 0.166 | 11.914 | 0.781 | 1.851 | 0.310 | 2.029 | 0.109 | 5.012 | 0.330 | 4.265 | 0.333 | ||||||
| HKU-Mars [19] | 2.695 | 1.208 | 2.399 | 1.956 | 1.814 | 0.706 | 2.578 | 1.086 | 2.527 | 0.246 | 14.996 | 3.386 | 3.941 | 1.261 | ||||||
| DVL [32] | 0.193 | 0.063 | 0.298 | 0.124 | 0.200 | 0.087 | 0.181 | 0.052 | 0.391 | 0.030 | 1.747 | 0.377 | 0.423 | 0.100 | ||||||
| MIAS-LCEC [4] | 0.190 | 0.050 | 0.291 | 0.111 | 0.198 | 0.072 | 0.177 | 0.046 | 0.363 | 0.024 | 0.749 | 0.118 | 0.298 | 0.061 | ||||||
| EdO-LCEC (Ours) | 0.168 | 0.044 | 0.293 | 0.105 | 0.184 | 0.057 | 0.183 | 0.044 | 0.338 | 0.027 | 0.474 | 0.104 | 0.255 | 0.055 | ||||||
In our experiments, we evaluate our proposed EdO-LCEC on four public datasets: KITTI odometry [30] (including 00-09 sequences), KITTI360 [45], nuScenes [46], and MIAS-LCEC [4] (including target-free datasets MIAS-LCEC-TF70 and MIAS-LCEC-TF360). Extensive comparisons with SoTA LCEC approaches and abundant ablation studies are conducted to comprehensively validate each component of the proposed algorithm. Following previous work [25, 4, 28, 32], we utilize the average magnitude of Euler angle error and the average magnitude of the translation error to quantify the calibration errors. Each axes of the Euler angle error and translation error are also provided to comprehensively demonstrate the calibration accuracy.
Notably, sequences in KITTI odometry, aside from 00, were included in the training datasets for the learning-based methods [25, 29, 27]. To ensure a fair comparison, we reproduced calibration results for both the left and right cameras on sequence 00 when the authors provided their code; otherwise, we used the reported results for the left camera from their papers. Since most learning-based methods lack APIs for custom data, our comparison with these methods is limited to the KITTI odometry 00 sequence. In nuScenes, the point clouds are captured by a 32-line spinning LiDAR (Velodyne HDL32E), which makes them extremely sparse for previous calibration approaches. To the best of our knowledge, few LCEC methods have been evaluated on this dataset. Following the official split, we use the scenes in v1.0-test of nuScenes to construct the evaluation data. For experiments on the MIAS-LCEC dataset, as the results for the compared methods [40, 19, 32, 4, 20] are reported in [4], we directly use the values presented in that paper.